OpenStudy (anonymous):

HELP! Solving functions? f(x)=1+√x-1 Solve each f(x): (1.) f(a+1) (2.) f(9) (3.) [f(x)]^2

OpenStudy (anonymous):

\[f(a+1)=1+\sqrt{a+1-1}=1+\sqrt{a}\]

OpenStudy (anonymous):

can you check my answers? 2) f(9)=1+2sqrt 2 3) 2(sqrt x-1) +x

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yep, they're both correct

OpenStudy (anonymous):

thanks could someone also check this: 2f(x)-1 i got this answer: 1+2 sqrt (x-1)

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

thanks a million

OpenStudy (anonymous):

can i start a new question and tag you in it to help?

OpenStudy (anonymous):

sure :)